OpenStudy (anonymous):

What is the volume of a cube with an area of 384cm2

OpenStudy (anonymous):

take the square root of what the area. so one of the sides= \[\sqrt{384}\]

OpenStudy (anonymous):

I get something like 19.596

OpenStudy (anonymous):

Now Volume = s^3

OpenStudy (anonymous):

so cube your answer.

OpenStudy (anonymous):

Volume = (\[\sqrt{384}^{3}\]

OpenStudy (anonymous):

that's your answer, I believe

OpenStudy (amistre64):

what kind of area?

OpenStudy (amistre64):

if its talking surface area, then 6*side area = 384 side area = 384/6 s^2 = 384/6 s = sqrt(384/6) volume is just s^3